# General quantum backflow in realistic wave packets

### v1

@@ Abstract
Introduction. Interference of matter waves—evidenced by the appearance of characteristic density fringes obtained through repeated position measurements of a particle at a given time—has been a cornerstone of quantum mechanics since its early days. Only relatively recently, however, has it been recognized that performing position measurements at multiple distinct times can reveal nonclassical aspects of probability transport even for a free nonrelativistic particle. Perhaps the most well-known example is quantum backflow [1, 2, 3, 4, 5], but other notable effects include quantum reentry [6], Tsirelson’s precession [7, 8, 9, 10, 11], and quantum advantage in various transportation tasks [12, 13, 14, 15]. Although these effects appeal to the very core of quantum mechanics—much like interference and diffraction—even quantum backflow has not yet been experimentally observed. [^1] Optical analogs of quantum backflow have been demonstrated for linear motion using classical beams [16, 17], and for rotational motion using both classical light [18] and single photons [19]. However, backflow of matter waves (i.e., for massive quantum particles) has not yet been observed.
@@
which holds for any $t_{0}<t_{1}<t_{2}$. [^2] See Supplemental Material for alternative derivation of Eqs. (1) and (2). For $P_{r}(t_{0})=1$ and $s\to\infty$, it reduces to the classical reentry condition discussed in Refs. [6, 30].
@@
Simple example. We now demonstrate that the classical bounds expressed in inequalities (1) and (2) can be violated by quantum particles. Consider a wave function consisting of a superposition of two Gaussian wave packets, $\Psi(x,t)=\sum_{n=1,2}c_{n}g_{n}(x,t)$. This type of superposition was previously analyzed in Ref. [23] for real coefficients $c_{n}$ and Gaussians $g_{n}(x,t)$ with identical widths, yielding quantum backflow of $0.0061$. Although strictly nonzero, the corresponding probability of negative momentum was found to be small, $\tilde{P}_{-}\approx 10^{-10}$, and was therefore neglected.
@@
Large backflow example. The previous illustration showed that general quantum backflow can exceed the magnitude observed in the standard scenario. Remarkably, general backflow can even significantly surpass the Bracken–Melloy constant, $c_{\text{BM}}=0.0384506$. Inspired by Ref. [24], we construct an explicit example using the momentum-space wave function $\tilde{\psi}(p)=\Theta(p)a_{1}(b_{1}-p)e^{-c_{1}^{2}p^{2}}+\Theta(-p)a_{2}(b_{2}-p)e^{-c_{2}^{2}p^{2}}$, where $a$’s, $b$’s, and $c$’s are constants, and $\Theta(\cdot)$ is the Heaviside step function. With a suitable choice of parameters, this state achieves a general backflow of $\Delta_{\mathrm{QB}}=0.0624$. There also exists a corresponding state exhibiting the same value of $\Delta_{\mathrm{RE}}$. Detailed derivations and parameter choices are provided in the SM.
@@
$\varphi(u)=\left(\frac{4\hbar m}{t_{2}-t_{1}}\right)^{\!\!1/4}\exp\!\left(-i\frac{t_{2}+t_{1}}{t_{2}-t_{1}}u^{2}\right)\tilde{\psi}\!\left(u\sqrt{\frac{4\hbar m}{t_{2}-t_{1}}}\right)$ (9)
@@
$\times$ $\psi\!\left(-u\sqrt{\frac{4\hbar\tau_{2}\tau_{1}}{m(\tau_{2}-\tau_{1})}}\right)$
@@
Supremum of $\Delta$. The maximal quantum violation of the classical inequalities (1) and (2) is obtained by optimizing $\Delta$ over all normalized functions $\varphi(u)$. Following Ref. [4], this is achieved by performing an unconstrained optimization of $\Delta+\lambda\int_{-\infty}^{+\infty}du\,|\varphi(u)|^{2}$, where $\lambda$ is a Lagrange multiplier. The resulting Euler-Lagrange equation reads
@@
Discussion. We conclude with several remarks on our results and their foundational implications. The central outcome of this work is a general characterization of quantum backflow and reentry for arbitrary wave functions, without imposing constraints on their momentum or position marginals.
@@
In the general backflow formulation, we considered the probability of locating the particle to the left of the origin. Naturally, the origin can be replaced by any point $b$ on the $x$ axis. Quantum backflow occurs whenever $\Delta_{\text{QB}}=\operatorname{Prob}(x<b,t_{2})-\operatorname{Prob}(x<b,t_{1})-\tilde{P}_{-}$ [c.f. Eq. (3)] becomes positive. Similarly, one can define quantum “forwardflow”, which occurs when $\Delta_{\text{QF}}=\operatorname{Prob}(x>f,t_{2})-\operatorname{Prob}(x>f,t_{1})-\tilde{P}_{+}$ is positive, with $f$ an arbitrary point and $\tilde{P}_{+}=1-\tilde{P}_{-}$ the probability of positive momentum. This raises a question: Can a quantum particle simultaneously exhibit backflow and forwardflow? In other words, does there exist a wave function for which $\Delta_{\text{QB}}>0$ and $\Delta_{\text{QF}}>0$ for some choice of parameters? The answer appears to be negative. It can be shown (see the SM for details) that $\Delta_{\mathrm{QB}}+\Delta_{\mathrm{QF}}\leq 0$ for all choices of parameters, revealing a trade-off and mutual incompatibility between these two effects.
@@
The Bracken–Melloy constant $c_{\text{BM}}$ characterizes maximal standard (positive-momentum) backflow and is independent of the particle’s mass, the backflow duration, and Planck’s constant. It was proposed to reflect the mathematical structure of quantum free-particle evolution [4]. The same properties hold for the general backflow constant $\sup\Delta$, given by Eq. (13), suggesting that quantum free evolution is characterized by at least two fundamental dimensionless constants. Understanding the relation between them presents a promising direction for future research.
@@
Finally, we comment on the foundational implications of quantum backflow and reentry. The classical limits, Eqs. (1) and (2), rely on two assumptions: (i) the existence of a joint probability density $f(x,p,t)$, and (ii) the validity of classical equations of motion for a free particle, $x_{t}=x_{0}+pt/m$ and $p_{t}=\text{const}$. Quantum backflow or reentry violates at least one of these assumptions and accordingly can be interpreted in different ways. Relaxing (i) while keeping (ii) leads to pseudo-probability frameworks, such as the Wigner function formalism, in which backflow emerges as negative probability transfer of particles in straight line motion [4, 23]. Conversely, retaining (i) and relaxing (ii) leads to Bohmian mechanics, where even free particles follow curved trajectories due to an effective nonlocal potential [32, 21].

### v2

# arXiv:2511.10155v2[quant-ph] 02 Mar 2026

@@ Abstract
Introduction. Interference of matter waves—evidenced by the appearance of characteristic density fringes obtained through repeated position measurements of a particle at a given time—has been a cornerstone of quantum mechanics since its early days. Only relatively recently, however, has it been recognized that performing position measurements at multiple distinct times can reveal nonclassical aspects of probability transport even for a free nonrelativistic particle. Perhaps the most well-known example is quantum backflow [1, 2, 3, 4, 5], but other notable effects include quantum reentry [6], Tsirelson’s precession [7, 8, 9, 10, 11, 12], and quantum advantage in various transportation tasks [13, 14, 15, 16]. These effects appeal to the very core of quantum mechanics, much like interference and diffraction. Yet even quantum backflow, which manifests as a classically impossible transfer of probability against the particle’s momentum, has not yet been experimentally observed. [^1] There has been significant experimental progress in studying optical analogs of quantum backflow, where the instantaneous local momentum of a light field is shown to point opposite to the local momentum of its constituent modes. Such effects have been demonstrated in linear motion using classical optical beams [17, 18], and in rotational motion using both classical light [19] and single photons [20]. However, because these experiments are effectively time-independent, they do not capture the probability dynamics that fundamentally define quantum backflow.
@@
which holds for any $t_{0}<t_{1}<t_{2}$. [^2] See Supplemental Material for an alternative derivation of Eqs. (1) and (2). For $P_{r}(t_{0})=1$ and $s\to\infty$, it reduces to the classical reentry condition discussed in Refs. [6, 31].
@@
Simple example. We now demonstrate that the classical bounds expressed in inequalities (1) and (2) can be violated by quantum particles. Consider a wave function consisting of a superposition of two Gaussian wave packets, $\Psi(x,t)=\sum_{n=1,2}c_{n}g_{n}(x,t)$, directly accessible in optical systems and Bose-Einstein condensates (BECs). This type of superposition was previously analyzed in Ref. [24] for real coefficients $c_{n}$ and Gaussians $g_{n}(x,t)$ with identical widths, yielding quantum backflow of $0.0061$. Although strictly nonzero, the corresponding probability of negative momentum was found to be small, $\tilde{P}_{-}\approx 10^{-10}$, and was therefore neglected.
@@
Large backflow example. The previous illustration showed that general quantum backflow can exceed the magnitude observed in the standard scenario. Remarkably, general backflow can even significantly surpass the Bracken–Melloy constant, $c_{\text{BM}}=0.0384506$. Inspired by Ref. [25], we construct an explicit example using the momentum-space wave function $\tilde{\psi}(p)=\Theta(p)a_{1}(b_{1}-p)e^{-c_{1}^{2}p^{2}}+\Theta(-p)a_{2}(b_{2}-p)e^{-c_{2}^{2}p^{2}}$, where $a$’s, $b$’s, and $c$’s are constants, and $\Theta(\cdot)$ is the Heaviside step function. Such superpositions are feasible in BEC platforms. With a suitable choice of parameters, this state achieves a general backflow of $\Delta_{\mathrm{QB}}=0.0624$. There also exists a corresponding state exhibiting the same value of $\Delta_{\mathrm{RE}}$. Detailed derivations and parameter choices are provided in the SM.
@@
$\varphi(u)=\left(\frac{4\hbar m}{t_{2}-t_{1}}\right)^{\!\!1/4}\exp\!\left(-i\frac{t_{2}+t_{1}}{t_{2}-t_{1}}u^{2}\right)\tilde{\psi}\!\left(u\sqrt{\frac{4\hbar m}{t_{2}-t_{1}}}\right)\,.$ (9)
@@
$\times$ $\psi\!\left(-u\sqrt{\frac{4\hbar\tau_{2}\tau_{1}}{m(\tau_{2}-\tau_{1})}}\right)\,$,
@@
Supremum of $\Delta$. The maximal quantum violation of the classical inequalities (1) and (2) is obtained by optimizing $\Delta$ over all normalized functions $\varphi(u)$. This is achieved by performing an unconstrained optimization of $\Delta+\lambda\int_{-\infty}^{+\infty}du\,|\varphi(u)|^{2}$, where $\lambda$ is a Lagrange multiplier [4]. The resulting Euler-Lagrange equation reads
@@
Discussion. We conclude with remarks on the foundational implications of our results. In the general backflow formulation, we considered the probability of locating the particle to the left of the origin. Naturally, the origin can be replaced by any point $b$ on the $x$ axis. Quantum backflow occurs whenever $\Delta_{\text{QB}}=\operatorname{Prob}(x<b,t_{2})-\operatorname{Prob}(x<b,t_{1})-\tilde{P}_{-}$ [c.f. Eq. (3)] becomes positive. Similarly, one can consider quantum “overflow”, recently introduced in Ref. [23], which occurs when $\Delta_{\text{QO}}=\operatorname{Prob}(x>f,t_{2})-\operatorname{Prob}(x>f,t_{1})-\tilde{P}_{+}$ becomes positive, with $f$ an arbitrary point and $\tilde{P}_{+}=1-\tilde{P}_{-}$ the probability of positive momentum. Since in the general formulation introduced here the particle can move in any direction, a natural question arises: can a particle simultaneously exhibit backflow and overflow? In other words, does there exist a wave function for which $\Delta_{\text{QB}}>0$ and $\Delta_{\text{QO}}>0$ for some choice of parameters? The answer is negative. It can be shown (see SM) that $\Delta_{\mathrm{QB}}+\Delta_{\mathrm{QO}}\leq 0$ for all parameters, revealing a trade-off and mutual incompatibility between these two effects. It is also worth noting that while quantum overflow can occur only over multiple disjoint time intervals in states with strictly positive momentum [23], it can already arise within a single interval for states with an arbitrary momentum distributions. In fact, quantum backflow and overflow are qualitatively equivalent phenomena, distinguished only by the direction of the probability transport.
@@
The Bracken–Melloy constant $c_{\mathrm{BM}}$ characterizes maximal standard (positive-momentum) backflow and is independent of the particle’s mass, the backflow duration, and Planck’s constant. It was proposed to reflect the mathematical structure of quantum free-particle evolution [4]. The same properties hold for the general backflow constant $\sup\Delta$, given by Eq. (13), suggesting that quantum free evolution is characterized by at least two fundamental dimensionless constants. Understanding the relation between them presents a promising direction for future research.
A more pragmatic follow-up question is how the strength of general backflow changes when the experimental evaluation of $\operatorname{Prob}(x<0,t)$ uses a smoothed spatial window rather than a sharp one. In the case of the standard positive-momentum formulation, such smoothing is known to lead to an underestimation of quantum backflow [24]. Another important aspect of a future experimental validation of general backflow is ensuring that any violation of inequalities (1) or (2) arises from quantum interference rather than from a failure of the dynamical assumptions (e.g., force-free motion) of the system under investigation. A theoretical framework for eliminating such a dynamical loophole in the context of the Tsirelson precession problem has recently been developed in Ref. [12].
@@
The classical limits, Eqs. (1) and (2), rely on two assumptions: (i) the existence of a joint probability density $f(x,p,t)$, and (ii) the validity of classical equations of motion for a free particle, $x_{t}=x_{0}+pt/m$ and $p_{t}=\text{const}$. Quantum backflow, reentry, and other nonclassical probability transport effects violate at least one of these assumptions. Relaxing (i) while keeping (ii) leads to pseudo-probability frameworks, such as the Wigner function formalism, in which backflow emerges as negative probability transfer of particles in straight line motion [4, 24]. Conversely, retaining (i) and relaxing (ii) leads to Bohmian mechanics, where even free particles follow curved trajectories due to an effective nonlocal potential [33, 22]. Yet, these trajectories have to comply with the Ehrenfest theorem. Moreover, one can demonstrate that any covariance measurements remain compatible with the straight line motion. Consequently, any signatures of backflow can only appear in the dynamics of third and higher moments of the position distribution. This demonstrates both the subtlety of the effect and the sensitivity of our methods to these higher-order changes.
We also note that the derivation of the inequalities (1) and (2) does not require strictly free-particle equations of motion. Rather, they hold for any continuous phase-space flow satisfying $\operatorname{sgn}(dx_{t}/dt)=\operatorname{sgn}(p_{t})$ and $\operatorname{sgn}(p_{t})=\text{const}$, where $\operatorname{sgn}(\cdot)$ denotes the sign function. These conditions require that the particle’s velocity be aligned with its momentum and that the direction of motion be preserved. This observation opens the possibility of investigating general quantum backflow and reentry beyond the free-particle setting.
@@
Acknowledgments. The authors thank Anat Daniel, Bohnishikha Ghosh, Bernard Gorzkowski, Thomas Juffmann, Radosław Łapkiewicz, Timon Zipfelmaier for insightful discussions.